RC Meeting, Saturday, October 23 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. Trippy beam came back Friday evening. It was possible to check the beam spot size and beam vs target alignment, but there was no significant production overnite. The beam current was briefly reduced to Halls A and C at one point to see if that would reduce trips.
  2. From the PD update this morning, a major source of problems was BLM trips near the now unpowered magnet in Arc 7. It sounds like the stray field on the adjacent arc had dramatically changed, so the thought is that the operators have to steer around that.
  3. As of 08:48, the shift crew reported that Hall C was back to 42 muA. But there is an unresolved issue of BLM trips seemingly dominated by a tail of the A beam going thru the C slit. See https://logbooks.jlab.org/entry/3930908 . MCC is delivering beam by reducing both A and C currents.
  4. Pass change for Hall A is possibly Monday.
  5. The accelerator will be off 09:30 Monday morning to prepare for the Secretary of Energy tour. Tours wrap up at 13:30.

Hall C Status

  1. We're taking beam of sketchy quality at reduced current, 30 muA, as of 10AM Saturday.
  2. I'm not updating the number of good pions until I have a better idea if any of this beam since Friday nite is useful.

Current Run Plan

  1. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  2. Q3 insulating vacuum spikes generating alarms. For example, https://logbooks.jlab.org/entry/3930175 .
  3. One HB LHe incident of note: https://logbooks.jlab.org/entry/3929152 . (possible clue to the origin?)
  4. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  5. Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
  6. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links